OpenStudy (swissgirl):

Prove For every real number A>0, there is a natural number B such that for all real numbers C>B, 1/(4C)

OpenStudy (kinggeorge):

Let's see here. Case 1: \(A\geq1\). If you take \(B=A\) then you have that \(4C\geq4A\). So you have that \[\frac{1}{4C}\leq\frac{1}{4A}\]Since \(A\geq1\), \(\frac{1}{A}\leq A\) with equality when \(A=1\), so \[\frac{1}{4C}<A\] Now we need a case 2 with \(A<1\).

OpenStudy (anonymous):

there is something wrong with your question ! you can't have for all real numbers 1/(4C)<A

OpenStudy (anonymous):

am I right KingGeorge ?!

OpenStudy (kinggeorge):

Case 2: \(A<1\). Now take \(B=\lceil \frac{1}{A}\rceil\). In this case, \(C>\frac{1}{A}\) so \[\frac{1}{C}<A\]Which implies \[\frac{1}{4C}<A\] I'm pretty sure this statement is in fact, true.

OpenStudy (swissgirl):

Yes it is true

OpenStudy (kinggeorge):

Note: In case 2, you should take \(B=\lceil \frac{1}{A}\rceil +1\) to ensure that \(C>\frac{1}{A}\).

OpenStudy (swissgirl):

okk gotcha still need to reread the proof to process it

OpenStudy (anonymous):

I'm sure it can't be true ! you can't choose the C !

OpenStudy (anonymous):

sorry ! I didn't understand this correctly :D ! for all real nnumers > B ...we have..... sorry...for wasting your time :D

OpenStudy (swissgirl):

Ya im like that too i read it too fast and totally misread it

OpenStudy (swissgirl):

Dont say u r like that KG cuz u r not

OpenStudy (kinggeorge):

It took me a few minutes to decide if this was true or false. It's certainly not intuitive.

OpenStudy (swissgirl):

Thanks for helpin me out :D

OpenStudy (kinggeorge):

No problem

OpenStudy (swissgirl):

umm King can i ask u a question?

OpenStudy (swissgirl):

Y did u have to split it up into 2 cases where A>1 and where A<1

OpenStudy (kinggeorge):

You might not have to. It was part of my process of determining if it was actually true or not. However, the argument for case 1 doesn't work for case 2, the argument for case 2 doesn't work for case 1. So if we didn't split it up into two cases, we would need an entirely different argument.

OpenStudy (swissgirl):

well lets say I wld say b=(1/(2a)). i mean it cld be niething

OpenStudy (kinggeorge):

If \(B=\lceil\frac{1}{4A}\rceil+1\), then \(C>\frac{1}{4A}\) so we know that \[4C>\frac{1}{A}\implies\frac{1}{4C}<A\]That seems to work.

OpenStudy (swissgirl):

okkkk yupppp

OpenStudy (kinggeorge):

If we choose \(A=10\) then \(B=2\), so \(C>2\). So \[\frac{1}{4C}<\frac{1}{8}<A\]So that seems to work.

OpenStudy (swissgirl):

okkkk gotcha thanks

OpenStudy (kinggeorge):

you're welcome.